# Evaluating infectious disease forecasts with allocation scoring rules

### 24 Feb 2024 v2

@@ B.1 The allocation score is proper
Appendix C Expected shortages
A key feature of loss functions for decision problems used to define quantiles and related scoring rules such as the CRPS and the WIS (see e.g., Gneiting (2011b), Jose and Winkler (2009), and Royset and Wets (2022), sections 1.C and 3.C), as well as the allocation loss function presented in this work, is the presence of a shortage: the amount $\max\{0,y-x\}$ by which a resource demand $y$ exceeds a supply decision variable $x$, which, for convenience, we write as $(y-x)_{+}$. In particular, a quantile at probability level $\alpha$ of a distribution $F$ on $\mathbb{R}^{1}$ (which we assume to have a well-defined density $f(x)$) is a Bayes act for the loss function
@@ Appendix C Expected shortages
Note that more care is required when $F$ does not have a density. We will also use this result below in deriving the Bayes act for the allocation loss.
Appendix D Allocation Bayes acts as vectors of marginal quantiles.
$\underset{0\leq x}{\mathrm{minimize}}\,\,\mathbb{E}_{F}[s_{A}(x,Y)]=\sum_{i=1}^{N}L\cdot\mathbb{E}_{F_{i}}[(Y_{i}-x_{i})_{+}]\text{ subject to }\,\sum_{i=1}^{N}x_{i}=K,$ (14)
In order for $x^{\star}\in\mathbb{R}^{N}_{+}$ to solve the AP it must be true that reallocating $\delta>0$ units of the resource from location $i$ to location $j$ will lead to a net increase in expected shortage — in other words, the reallocation increases the expected shortage in location $i$ at least as much as it decreases the expected shortage in location $j$:
$\mathbb{E}_{F_{i}}[(Y_{i}-(x^{\star}_{i}-\delta))_{+}]-\mathbb{E}_{F_{i}}[(Y_{i}-x^{\star}_{i})_{+}]\text{ (increase in $i$) }$
$\qquad\geq\mathbb{E}_{F_{j}}[(Y_{j}-x^{\star}_{j})_{+}]-\mathbb{E}_{F_{j}}[(Y_{j}-(x^{\star}_{j}+\delta))_{+}]\text{ (decrease in $j$) }.$
@@
$1-F_{i}(x^{\star}_{i})$ $=-\frac{d}{dx_{i}}\mathbb{E}_{F_{i}}[(Y_{i}-x^{\star}_{i})_{+}]$
$=\lim_{\delta\searrow 0}\frac{1}{\delta}\left\{\mathbb{E}_{F_{i}}[(Y_{i}-(x^{\star}_{i}-\delta))_{+}]-\mathbb{E}_{F_{i}}[(Y_{i}-x^{\star}_{i})_{+}]\right\}\text{ (increase in $i$) }$
$\geq\lim_{\delta\searrow 0}\frac{1}{\delta}\left\{\mathbb{E}_{F_{j}}[(Y_{j}-x^{\star}_{j})_{+}]-\mathbb{E}_{F_{j}}[(Y_{j}-(x^{\star}_{j}+\delta))_{+}]\right\}\text{ (decrease in $j$) }$
$=-\frac{d{}}{d{x_{j}}}\mathbb{E}_{F_{j}}[(Y_{j}-x^{\star}_{j})_{+}]=1-F_{j}(x^{\star}_{j})$ (15)
Note that negative derivatives appear because our optimality condition addresses how a decrease in resources will increase the expected shortage in $i$ and vice versa in $j$. Since (15) also holds with $i$ and $j$ reversed, a number $\lambda$ (a Lagrange multiplier) exists such that $L(1-F_{k}(x^{\star}_{k}))=\lambda$ for all $k\in 1,\ldots,N$. (We scale by $L$ to facilitate possible future interpretations of $\lambda$ in terms of the partial derivatives of $\mathbb{E}_{F}[s_{A}(x,Y)]$.) That is, $x^{\star}_{k}$ is a quantile $q_{\tau,F_{k}}$ for $\tau=1-\lambda/L$. The value of $\tau$ is then determined by the constraint equation
@@
Appendix E Numerical computation of allocation Bayes acts
@@
Appendix F Computing allocations from finite quantile forecast representations
@@
To address this situation and support the numerical allocation algorithm outlined in section E, we need a mechanism to approximate the full cumulative distribution functions $F_{i}$, $i=1,\ldots,N$ based on the provided quantiles. We have developed functionality for this purpose in the distfromq package for R (Ray and Gerding, 2023). This functionality represents a distribution as a mixture of discrete and continuous parts, and it works in two steps:
@@
Appendix G Propriety of parametric approximation
@@
G.1 Propriety when scoring methods are announced prospectively
@@
It follows from the discussion in section B.1 that this is a proper scoring rule for $F$. Although the full forecast distribution $F$ is not available in the forecast submission, the score $S_{R}(F,y)$ can be calculated from the reported parameter values as long as the forecaster submits the optimal parameters $\theta^{\star}(F)$.
@@
G.2 Impropriety of post hoc allocation scoring with quantile forecasts

### 04 Mar 2024 v3

@@ A.1 The allocation score is proper
Appendix B Allocation Bayes acts as vectors of marginal quantiles.
$\underset{0\leq x}{\mathrm{minimize}}\,\,\mathbb{E}_{F}[s_{A}(x,Y)]=\sum_{i=1}^{N}L\cdot\mathbb{E}_{F_{i}}[\max(0,Y_{i}-x_{i})]\text{ subject to }\,\sum_{i=1}^{N}x_{i}=K,$ (12)
We begin by noting that a key feature of each term of the loss function $s_{A}(x,Y)$ defining the AP is the presence of a shortage: an amount $\max\{0,y-x\}$ by which a resource demand $y$ exceeds a supply decision variable $x$, which, for convenience, we write as $(y-x)_{+}$. This is a feature shared with decision problems used to define quantiles and related scoring rules such as the CRPS and the WIS (see e.g., Gneiting (2011b), Jose and Winkler (2009), and Royset and Wets (2022), sections 1.C and 3.C). In particular, a quantile at probability level $\alpha$ of a distribution $F$ on $\mathbb{R}$ (which we assume to have a well-defined density $f(x)$) is a Bayes act for the loss function
@@
Note that more care is required when $F$ does not have a density.
Returning to the AP (12), notice that in order for $x^{\star}\in\mathbb{R}^{N}_{+}$ to be a Bayes act it must be true that reallocating $\delta>0$ units of the resource from location $i$ to location $j$ will lead to a net increase in expected shortage — in other words, the reallocation increases the expected shortage in location $i$ at least as much as it decreases the expected shortage in location $j$:
$\mathbb{E}_{F_{i}}[(Y_{i}-(x^{\star}_{i}-\delta))_{+}]-\mathbb{E}_{F_{i}}[(Y_{i}-x^{\star}_{i})_{+}]\text{ (detriment in $i$) }$
$\qquad\geq\mathbb{E}_{F_{j}}[(Y_{j}-x^{\star}_{j})_{+}]-\mathbb{E}_{F_{j}}[(Y_{j}-(x^{\star}_{j}+\delta))_{+}]\text{ (benefit in $j$) }.$
@@
$1-F_{i}(x^{\star}_{i})$ $=-\frac{d}{dx_{i}}\mathbb{E}_{F_{i}}[(Y_{i}-x^{\star}_{i})_{+}]\text{ (rate of detriment in $i$) }$
$=\lim_{\delta\searrow 0}\frac{1}{\delta}\left\{\mathbb{E}_{F_{i}}[(Y_{i}-(x^{\star}_{i}-\delta))_{+}]-\mathbb{E}_{F_{i}}[(Y_{i}-x^{\star}_{i})_{+}]\right\}$
$\geq\lim_{\delta\searrow 0}\frac{1}{\delta}\left\{\mathbb{E}_{F_{j}}[(Y_{j}-x^{\star}_{j})_{+}]-\mathbb{E}_{F_{j}}[(Y_{j}-(x^{\star}_{j}+\delta))_{+}]\right\}$
$=-\frac{d{}}{d{x_{j}}}\mathbb{E}_{F_{j}}[(Y_{j}-x^{\star}_{j})_{+}]=1-F_{j}(x^{\star}_{j})\text{ (rate of benefit in $j$).}$. (15)
(Negative derivatives appear here because our optimality condition addresses how a decrease in resources will increase the expected shortage in $i$ and vice versa in $j$.)
Since (15) also holds with $i$ and $j$ reversed, a number $\lambda$ (a Lagrange multiplier) exists such that $L(1-F_{k}(x^{\star}_{k}))=\lambda$ for all $k\in 1,\ldots,N$. (We scale by $L$ to facilitate possible future interpretations of $\lambda$ in terms of the partial derivatives of $\mathbb{E}_{F}[s_{A}(x,Y)]$.) That is, $x^{\star}_{k}$ is a quantile $q_{\tau,F_{k}}$ for $\tau=1-\lambda/L$. The value of $\tau$ is then determined by the constraint equation
@@
Appendix C Numerical computation of allocation Bayes acts
@@
Appendix D Computing allocations from finite quantile forecast representations
@@
To address this situation and support the numerical allocation algorithm outlined in section C, we need a mechanism to approximate the full cumulative distribution functions $F_{i}$, $i=1,\ldots,N$ based on the provided quantiles. We have developed functionality for this purpose in the distfromq package for R (Ray and Gerding, 2023). This functionality represents a distribution as a mixture of discrete and continuous parts, and it works in two steps:
@@
Appendix E Propriety of parametric approximation
@@
E.1 Propriety when scoring methods are announced prospectively
@@
It follows from the discussion in section A.1 that this is a proper scoring rule for $F$. Although the full forecast distribution $F$ is not available in the forecast submission, the score $S_{R}(F,y)$ can be calculated from the reported parameter values as long as the forecaster submits the optimal parameters $\theta^{\star}(F)$.
@@
E.2 Impropriety of post hoc allocation scoring with quantile forecasts